OpenStudy (anonymous):

using the properties of logs to verify, how do I get -(3 ln 2 + ln 3) from -ln 24

OpenStudy (anonymous):

2 points and a medal

OpenStudy (anonymous):

Thank you Xeph

OpenStudy (anonymous):

Starting with -ln 24 and trying to make it equal to the other side\[-\ln 24=-\ln(3\times8)\]\[=-(\ln3+\ln8)\]\[=-(\ln3+\ln2^3)\]\[=-(\ln3+3\ln2)\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem. Do you understand the logarithm laws I used?

OpenStudy (anonymous):

Kind of, but why did you multiply 3 and 8? I know to get 24, but can you always do that?

OpenStudy (anonymous):

I have to go to bed soon... are you always available around 6:30 - 8 almost every night on here?

OpenStudy (anonymous):

It might not be as obvious in other problems, but since I was trying to make it equal to the other side and i saw ln3 was one of the elements so dividing 3 out of 24 would likely be a good first step

OpenStudy (anonymous):

nah i have inconsistent times, but i'm on fairly often; i'm heading to bed now too

OpenStudy (anonymous):

Good night!

OpenStudy (anonymous):

you too

OpenStudy (anonymous):

Go Chicago Bulls!